A day on a distant planet observed orbiting a nearby star is 52.1 hours. Also, a year on the planet lasts 38.7 Earth days.* Calculate the average angular speed of the planet about its own axis of rotation in radians per second, with the second as measured on Earth.

Calculate the average angular speed of the planet as it travels around its neighboring star, with the second as measured on Earth.
Answer #1
a) w = 2pi/(52.1*3600) = 3.349960140329345*10^-5 rad/s

b) w= 2pi/(38.7*24*3600) = 1.8791227954131365*10^-6 rad/s